OpenStudy (anonymous):

x - rational number, y - not rational number, How to prove (formal) that x + y - not rational number

OpenStudy (anonymous):

are you saying x is irrational and y is irrational therefore x + y is irrational? you cannot prove this because it is not true

OpenStudy (anonymous):

for example \[\sqrt{2}\] is irrational, as is \[3-\sqrt{2}\] but their sum is \[\sqrt{2}+3-\sqrt{2}=3\]which is certainly rational

OpenStudy (anonymous):

But x is rational

OpenStudy (anonymous):

ooooooooh sorry i read incorrectly. sum of rational and irrational is irrational,

OpenStudy (anonymous):

I think i should show that x+y cannot be presented as a / b (where a,b - integers and b is not 0)

OpenStudy (anonymous):

ok so try this. if x is rational and y is irrational, let us assume for a moment that \[x+y\] is rational and derive a contradiction. since rational numbers are closed under subtraction , that means \[x+y-x=y \] is rational, but you are assuming that y is not rational, so you have a contradiction.

OpenStudy (anonymous):

a is a rational number \[a = \frac{m}{n}\] b is a irrational number Suppose that a + b is rational \[\frac{m}{n} +b = \frac{o}{p}\] \[b = \frac{o}{p} -\frac{m}{n}\] Since a rational number minus a rational number must equal to a rational number. Therefore a + b must be irrational

OpenStudy (anonymous):

Thanks a lot, moneybird and satelite73 ! I think that moneybird answer is better for me since it's based on a/b definition of rational numbers.